OpenStudy (anonymous):

How would you factorise x^2-y^2-2x-2x-6y-8.......??

OpenStudy (lgbasallote):

he was asking if that was the question lol

OpenStudy (cherylim23):

@moongazer, you can also factorise expressions, not only equations. x^2-y^2-2x-2x-6y-8 = x^2 -4x - y^2 - 6y-8 Complete the square... = (x-2)^2 -4 - [(y+3)^2 - 9] - 8 = (x-2)^2 - (y+3)^2 - 4 + 9 - 8 = (x-2)^2 - (y+3)^2 -3

OpenStudy (moongazer):

@cherylim23 That's what I am going to do. I just taught you can't use completing the square if it is not an equation. :)

OpenStudy (cherylim23):

Haha of course you can